a piano, mass unknown, is being hoisted up by a crane to the second floor of a building when the rope snaps. The piano drops 14 meters. What speed does it have just as it hits the ground? please explain why a certain formula is used
Answer #1

v^2 = u^2 + 2gh

here u = 0

v = sqrt(2gh)

v = 16.57 m/s

here I used this equation because we want to know the final speed , height and initial speed are known

so I used motions third equation